0r2 points Human visual wetting, knee visibility, volds) and even the degree to which a joint possesses one or more of these defects. Consequently, even highly trained inspectors can disagree on the disposition of a particular joint. In one batch of 10,000 joints, inspector A found 742 that were judged defective, inspector B found 732 such joints, and 1351 of the joints were judged defective by at least one of the inspectors. Suppose that one of the 10,000 jolints is randomly selected of solder joints on prnted circuit boards can be very subjective. Part of the problem stems from the numerous types of solder defects (e.g.. pad non (a) What is the probability that the selected joint was judged to be defective by neither of the two inspectors? (Enter your answer to four decimal places.) (b) What is the probability that the selected joint was judged to be defective by Inspector B but not by inspector A? (Enter your answer to four decimal places.) Activate Windows
Answer #1

P(A) = 742/10000 = 0.0742, P(B) = 732/10000 = 0.0732, P(At least one of A and B) = 1351/10000 = 0.1351

(a) P(Neither A nor B) = 1 - P(At least one of A and B) = 1 - 0.1351 = 0.8649

(b) n(A) = 742, n(B) = 732, n(A or B) = 1351

n(A and B) = n(A) + n(B) – n(A or B) = 742 + 732 – 1351 = 123

n(B but not A) = n(B) – n(A and B) = 732 – 123 = 609

P(B but not A) = 609/10000 = 0.0609